OpenStudy (anonymous):

okay can someone walk me through this, i dont have an example in my book for finding marginal demand. D(p) =(6p+250) / (8p+19), what is the marginal demand D'(p)

OpenStudy (e.mccormick):

Notice that it is D(p) given and D'(p) being asked for?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

And the prime means....

OpenStudy (anonymous):

its a derivative

OpenStudy (anonymous):

or i have to find the derivative

OpenStudy (e.mccormick):

D'(p) is a derivative. D(p) is not. So to get D'(p) from D(p).... yah, and they really expect you to figure that out just from one little mark.

OpenStudy (e.mccormick):

So quotent rule and there you go.

OpenStudy (anonymous):

okay i have the quotient rule. i'll give it a shot. think you can double check my answer for me in a minute?

OpenStudy (e.mccormick):

ok.

OpenStudy (anonymous):

2114 / [8p+19] squared ???

OpenStudy (e.mccormick):

2114? Fergot to distribute your -

OpenStudy (anonymous):

not seeing what you're asking. followed book example exactly

OpenStudy (e.mccormick):

\[\frac{6(8p+19)-8(6p+250)}{(8p+19)^2}\implies \frac{48p+144-48p-2000}{(8p+19)^2}\]

OpenStudy (e.mccormick):

Algebra mistake, the - distributes. 90% of calculus is algebra and trig and even arithmetic. That is where we all make the misakes.

OpenStudy (anonymous):

6x19 is 114

OpenStudy (e.mccormick):

Yes, but what is: \(-8\times 250\)?

OpenStudy (anonymous):

oh i see. you know ive missed questions on every test because i fail to account for negative numbers total mental block

OpenStudy (anonymous):

so the numerator is 1886

OpenStudy (e.mccormick):

Hehe. Yah, the minus sign is kind of like pants. Emberassing when dropped at the wrong time.

OpenStudy (e.mccormick):

\[\frac{-1886}{(8p+19)^2}\]

OpenStudy (anonymous):

thanks much for your help. I'm feeling alot more confident about m understanding of this stuff today. Hopefully it will stick.

OpenStudy (anonymous):

have a good one

OpenStudy (e.mccormick):

Don't have too much fun!